The measure of one angle is twice the measure of its complement, find the measure of the angles. Set up equation and solve
Answer by ikleyn(52792)

.

Let x be the angle measure, in degrees;

then the measure of its complement is 90-x degrees.


According the condition,


    x = 2*(90-x).


Simplify and solve to find x


    x = 180 - 2x,

    x + 2x = 180

    3x     = 180

     x     = 180/3 = 60 degrees.


ANSWER.  The angle is 60 degrees.  The complementary angle is 90-60 = 30 degrees.

Solved.

-------------

To learn about complementary/supplementary angles,  read the lessons
    - Angles basics
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
in this site.